# Forming the local starburst galaxy Haro 11 through hydrodynamical merger simulations

### v1

@@ 5.2 Varying velocity
We briefly overview the effects of varying the initial velocities, as velocity is connected to characteristics of the tidal tail and the stellar populations (the interaction times are connected to the SFR bursts). Generally, increasing the speed of the progenitors extended the time before the second encounter, giving the tidal tail more time to extend further. However, it also leads to a very brief interaction time which results in a weak tidal tail. In contrast, lower velocities lead to a quicker merger and shorter tidal tail, but the gas within the tail is denser.
@@ 5.3 Elliptical vs disc
A formation scenario proposed mentioned in Adamo et al. (2010) is that the progenitor galaxies of Haro 11 could be a low-mass, evolved galaxy and a gas rich (disc) galaxy, due to the irregular gas morphology around the knots and stellar age distribution (see discussion in Section 2). We briefly explored a setup where the retrograde disc was replaced by a gas-poor, dwarf elliptical galaxy. For this test, we kept the total mass of the systems the same, but the gas and stars was re-distributed between the galaxies to replace the prograde disc with a sphere [^3] The stellar distribution followed a Hernquist profile with similar scale radius as the original disc. of stars with no gas. An image of the tidal tail gas density is presented in Figure 13, which shows that this approach also exhibits only one tidal tail; as to be expected when there is no disc material to be stripped. Additionally, the extended spherical shape cause some additional tidal substructures to appear. We did not analyse the inner morphology.
The SFR of this scenario is lower leading up to the first encounter, after which is is comparable to other test simulations with discs. Although, it does not reach the same peak in SFR nor exhibit any distinct bump during the first encounter, compared to the fiducial simulation in Figure 7. This does not necessarily eliminate this specific merger scenario, as other changes, such as a higher gas fraction in the prograde galaxy, could be made to boost the SFR.
@@ 5.4 Density of the CGM
Additionally, we performed a few tests varying the circumgalactic density. The fiducial value $\rho_{\rm CGM}=10^{-4}\,{\rm m_{H}\,cm}^{-3}$ is calculated by spreading some of the gas mass in a 603 kpc3 cube, i.e. the size of the largest of the two dark matter halos. This gas mass was taken as the baryonic mass missing from the galaxy mass to reach the cosmic baryon fraction. In Figure 14 we show an example of how the increased density affects the tidal tail structure. The tail can be seen to go from being one coherent structure at low density and becomes clumpy towards higher density. Comparably, observations of the HI tail show that the tail is clumpy in nature (Le Reste et al., 2024), which could be due to interactions with a dens halo of gas or clumpiness within disc before or during stripping. Notably, a more gaseous halo would also affect the shock structure of ejected gas and could be compared with observed ionisation structures, but due to these examples having no stellar feedback there is no evident difference in outflow shocks.
@@
Figure 14: Density maps of the tidal tail in low-resolution simulations of varying circumgalatic density. From left to right, the densities are: ${\rm log}_{10}\,\rho_{\rm CGM}=(-6,-5,-4)\,{\rm m_{H}\,cm}^{-3}$.
@@ 6 Discussion
6.1 Caveats and alternative merger scenarios
We have in the previous sections demonstrated that our simulation setup of two disc galaxies undergoing a merger is able to reproduce the key features of the morphology, kinematics, SFR, and stellar population of Haro 11. Finding a model that is able to reproduce several different features of a specific galaxy is not trivial and is encouraging for the reliability of the presented merger scenario. However, there are caveats to our results: the model is not accurate enough for detailed comparison and there are degeneracies between different setups. Here, we connect back to the observational data to address the strength and weaknesses of our simulation model.
The simulations in this work show that the presence of a tidal tail, as reported by Le Reste et al. (2024), are in agreement with the scenario that Haro 11 is currently undergoing a merger. This is further confirmed by new observational data presented in this work of a stellar tidal tail, which is the first time the stellar component has been captured. However, because the Haro 11 tidal tail is observed as a 2D projection and does not show any curvature or other distinct morphological features, it is difficult to model the tidal tail morphology in detail, which limits the constraint to the length and width of the tail. This puts more emphasis on also matching the kinematics and other properties, such as the relative HI gas fraction. We find that the properties of the observed tail match with our fiducial Haro 11 simulation, and since the properties of the tidal tail are closely connected to the orbital conditions, this is a good indication that the orbit of the fiducial simulation are correct. Another sign of this is our discovery that a specific set of parameters are required for the second tidal tail to be undetectable.
The inner morphology of Haro 11 has a few distinct features that we are able to match: the three stellar knots, the disc arm of gas and stars (’ear’), the ’bent’ shape of the galaxy, in particular knot A and B being covered by gas. Fitting the exact positions of the knots (relative each other) and the orientation of the disc arm requires detailed tweaking of orbital parameters. Additionally, the relative position of the knots is closely connected to the viewing angle and the chosen time output. We note that by varying these, the position of these morphological features can be improved. However, changing the angles was found to lead to a severely worse match with other features; primarily the length, position, and kinematics of the tidal tail.
@@ 6.1 Caveats and alternative merger scenarios
From our mock observations in Figure 1, knot A is not evidently visible. This is likely due to the amount of dust and gas within the arm, obscuring the knot. Instead, we infer its presence from the stellar mass distribution, young stellar ages, and the placement of the disc arm. Additionally, the amount of young stars in knot A could affects its visibility. This depends on the specifics of small-scale physics such as the star formation efficiency and gas clumping, but also the gas fractions and timing of observations.
The final gas and stellar mass between simulation and observations are in agreement, but this is neither surprising nor a viable constraint as the galaxy are modelled to have similar masses. However, it is reassuring that no additional effects, such as stellar feedback and tidal stripping, have significantly altered this match. Furthermore, the comparison between gas phases is uncertain, but it is encouraging that there is a broad agreement in the distribution. Although, there is a significant amount ($1.90\times 10^{9}{\rm\,M_{\odot}}$) of hot/warm ionised gas within the halo of our galaxies, which is too diffuse to be observed and therefore presented separately from the ionised gas mass.
@@
The SFR of our chosen output is slightly higher than the observed SFR ($\sim 40{\rm\,M_{\odot}\,yr^{-1}}$ compared to $\lesssim 30{\rm\,M_{\odot}\,yr^{-1}}$). This increases to $\gtrsim 100{\rm\,M_{\odot}\,yr^{-1}}$ within the time it takes the galaxies to merge and the morphology to become unrecognisable. Even if we chose this later output, the factor of two difference in the SFR is well within the uncertainties of small-scale physics within the simulation that regulate star formation, e.g. the adopted star formation efficiency, stellar feedback, among other factors. We found that the SFR was stable to small tweaks to the orbit, with exception of close interactions. However, the SFR is highly sensitive to the chosen time output we choose to observe, due to the fact that the Haro 11 model is currently undergoing a powerful starburst which can vary from a few ${\rm\,M_{\odot}\,yr^{-1}}$ to $100{\rm\,M_{\odot}\,yr^{-1}}$ within $\sim 10$ Myr. Thus, the current SFR sets a relatively strong constraint for the time at which we observe the galaxy.
From exploring the orbital parameter space of this merger, we recognise there are numerous ways to match the features of Haro 11 and, thus, many possible formation scenarios. However, to find a set of parameters that produce several key aspects of the galaxy is not trivial. This is the first time Haro 11 has ever been simulated with hydrodynamical simulations, and that our simulations match several observed Haro 11 features is thus highly encouraging. Furthermore, it validates the hypothesis that Haro 11 is produced through a merger, and offers a specific possible merger setup. Future work, applying more precise constraints and comparing other aspects of the galaxy, would be helpful to reduce the available parameter space.
@@ 7 Conclusions
    • In order for the tidal tail to expand to the observed length of 40 kpc, our simulations require $200-250$ Myr between the first and second interaction.
@@
    • The galaxy centres will merge and become indistinguishable within $10-20$ Myr, creating another brief starburst period.
@@ Appendix A HI velocity map of the absorption
[…]
Figure 15: MEERKAT observations of the 21-cm line in absorption (Le Reste et al., 2024).

### v2

@@ 5.2 Varying velocity
We briefly review the effects of varying the initial velocities, as velocity is connected to characteristics of the tidal tail and the stellar populations (the interaction times are connected to the SFR bursts). Generally, increasing the speed of the progenitors extended the time before the second encounter, giving the tidal tail more time to extend further. However, it also leads to a very brief interaction time which results in a weak tidal tail. In contrast, lower velocities lead to a quicker merger and shorter tidal tail, but the gas within the tail is denser.
@@ 5.3 Elliptical vs disc
A formation scenario proposed mentioned in Adamo et al. (2010) is that the progenitor galaxies of Haro 11 could be a low-mass, evolved galaxy and a gas rich (disc) galaxy, due to the irregular gas morphology around the knots and stellar age distribution (see discussion in Section 2). We briefly explored a setup where the retrograde disc was replaced by a gas-poor, dwarf elliptical galaxy. For this test, we kept the total mass of the systems the same, but the gas and stars were redistributed between the galaxies to replace the prograde disc with a sphere [^2] The stellar distribution followed a Hernquist profile with similar scale radius as the original disc. of stars with no gas. An image of the tidal tail gas density is presented in Figure 13, which shows that this approach also exhibits only one tidal tail; as to be expected when there is no disc material to be stripped. Additionally, the extended spherical shape causes some additional tidal substructures to appear. We did not analyse the inner morphology.
The SFR of this scenario is lower leading up to the first encounter, after which it is comparable to other test simulations with discs. Although, it does not reach the same peak in SFR nor exhibit any distinct bump during the first encounter, compared to the fiducial simulation in Figure 7. This does not necessarily rule out this specific merger scenario, as other changes, such as a higher gas fraction in the prograde galaxy, could be made to boost the SFR.
@@ 5.4 Density of the CGM
Additionally, we performed a few tests varying the circumgalactic density. The fiducial value $\rho_{\rm CGM}=10^{-4}\,{\rm m_{H}\,cm}^{-3}$ is calculated by spreading some of the gas mass in a 603 kpc3 cube, i.e. the size of the largest of the two dark matter halos. This gas mass was taken as the baryonic mass missing from the galaxy mass to reach the cosmic baryon fraction. In Figure 14 we show an example of how the increased density affects the tidal tail structure. The tail can be seen to go from being one coherent structure at low density and becomes clumpy towards higher density. Comparably, observations of the HI tail show that the tail is clumpy in nature (Le Reste et al., 2024), which could be due to interactions with a dense halo of gas or clumpiness within disc before or during stripping. Notably, a more gaseous halo would also affect the shock structure of ejected gas and could be compared with observed ionisation structures, but due to these examples having no stellar feedback there is no evident difference in outflow shocks.
@@
Figure 14: Density maps of the tidal tail in low-resolution simulations of varying circumgalactic density. From left to right, the densities are: ${\rm log}_{10}\,\rho_{\rm CGM}=(-6,-5,-4)\,{\rm m_{H}\,cm}^{-3}$.
@@ 6 Discussion
6.1 Caveats and parameter variation
We have in the previous sections demonstrated that our simulation setup of two disc galaxies undergoing a merger is able to reproduce the key features of the morphology, kinematics, SFR, and stellar population of Haro 11. Finding a model that is able to reproduce several different features of a specific galaxy is not trivial and is encouraging for the reliability of the presented merger scenario. However, there are caveats to our results: the model is not accurate enough for detailed comparison and different setups can yield similar results (e.g. morphologies). Thus, it is difficult to attain a definitively best set of parameters. Here, we connect back to the observational data to address the strength and weaknesses of our simulation model.
6.1.1 Tidal tail
The simulation in this work show that the presence of a single tidal tail, as reported by Le Reste et al. (2024), is in agreement with the scenario that Haro 11 is currently undergoing a merger. This is further confirmed by new observational data presented in this work of a stellar tidal tail, which is the first time the stellar component has been captured. However, because the Haro 11 tidal tail is observed as a 2D projection and does not show any curvature or other distinct morphological features, it is difficult to model the tidal tail morphology in detail, which limits the constraint to the length and width of the tail. This puts more emphasis on also matching the kinematics and other properties, such as the relative HI gas fraction. Overall, we find that the properties of the observed tail match well with our fiducial Haro 11 simulation, and since the properties of the tidal tail are closely connected to the orbital conditions, this is a good indication that the orbit of the fiducial simulation is correct. Another good indication is our discovery that a specific set of parameters are required for the second tidal tail to not be present/detectable, which limits the allowed setups.
[…]
6.1.2 Inner morphology & kinematics
The inner morphology of Haro 11 has a few distinct features that we are able to match: the three stellar knots, the disc arm of gas and stars (’ear’), the ’bent’ shape of the galaxy, in particular knot A and B being covered by gas. Fitting the exact positions of the knots (relative each other) and the orientation of the disc arm requires detailed tweaking of orbital parameters. Additionally, the relative position of the knots is closely connected to the viewing angle and the chosen time output. We note that by varying these, the position of these morphological features can be improved. However, changing the angles was found to lead to a worse match with other features; primarily the length, position, and kinematics of the tidal tail.
@@ 6.1.2 Inner morphology & kinematics
From our mock observations in Figure 1, knot A is not visible. Instead, we infer its presence from the stellar mass distribution, young stellar ages, and the placement of the disc arm. In agreement with observations, there is a presence of dust and gas around knot A, but our simulation show more dust along the line of sight which obscures the knot. Additionally, the amount and density of young, luminous stars in knot A could affects its visibility. The formation of luminous clusters depends on the specifics of small-scale physics, such as the star formation efficiency and gas clumping, as well as the initial conditions, e.g. the gas fraction and disc properties. Additionally, the morphology depends on the timing and viewing angle for the analysis.
@@
6.1.3 Stellar and gas properties
The SFR of our chosen output is slightly higher than the observed SFR ($\sim 40{\rm\,M_{\odot}\,yr^{-1}}$ compared to $\lesssim 30{\rm\,M_{\odot}\,yr^{-1}}$). This increases to $\gtrsim 100{\rm\,M_{\odot}\,yr^{-1}}$ within the time it takes the galaxies to merge and the morphology to become unrecognisable. Even if we chose this later output, the factor of two difference in the SFR is well within the uncertainties of small-scale physics within the simulation that regulate star formation, e.g. the adopted star formation efficiency and stellar feedback. We found that the SFR was stable to small tweaks to the orbit, with the exception of close interactions. However, the SFR is highly sensitive to the time output we choose to analyse the galaxy, due to the fact that the Haro 11 model is currently undergoing a powerful starburst which can vary from a few $\rm\,M_{\odot}\,yr^{-1}$} to $100{\rm\,M_{\odot}\,yr^{-1}}$ within $\sim 10$ Myr. Thus, the current SFR sets a relatively strong constraint for the time at which we observe the galaxy.
The stellar clusters in our simulation show similar findings as in observations: knot B and C are more massive and contain older stars, compared to knot A. However, our simulation suggest the formation of more massive clusters, which would be readily observed if real. These massive clusters are partially a side-effect of how the post-processing clustering algorithm works, as we chose parameters that enhanced clustering to highlight low mass clusters. Furthermore, we are not fully capturing all stellar cluster (in particular the least massive) due to the limited spatial and mass resolution in our simulations.
The final gas and stellar mass between simulation and observations are in agreement, but this is neither surprising nor a viable constraint as the galaxy is modelled to have similar masses. However, it is reassuring that no additional effects, such as stellar feedback and tidal stripping, have significantly altered this match. Furthermore, the comparison between gas phases is uncertain, but it is encouraging that there is a broad agreement in the distribution. Although, there is a significant amount ($1.90\times 10^{9}{\rm\,M_{\odot}}$) of hot/warm ionised gas within the halo of our galaxies, which is too diffuse to be observed and therefore presented separately from the ionised gas mass.
[…]
@@ 7 Conclusions
    • In order for the tidal tail to expand to the observed length of 40 kpc, our simulations require 200250200-250 Myr between the first and second interaction.
@@
    • The galaxy centres will merge and become indistinguishable within 102010-20 Myr, creating another brief starburst period.
@@ Appendix A HI velocity map of the absorption
The velocity map of neutral HI emission presented in Figure 3 shows emission in both the inner parts and tidal tail. Meanwhile, the absorption presented here in Figure 15 shows significant absorption only in the inner parts of the galaxy. Furthermore, it indicates that there is a large outflow of neutral gas at velocities $\sim 50$$\rm{km\,s^{-1}}$} towards us.
Figure 15: Moment-1 (velocity centroid) map of the 21-cm line in absorption from MeerKAT observations (Le Reste et al., 2024).
Appendix B Optical observations of stellar tidal tail
The detection of the stellar tidal tail in Haro 11, as discussed and shown in Section 4.1, was done with VLT/FORS (program 113.26LV). Observations were done for a net integration time of 6240s in the B-band (filter b_HIGH+113 split in 8 exposures of 780s) and 5760s in I (I-BESS+77, split in 24 exposures of 240s). The shorter exposure time in I-band was chosen to ensure that the sky background did not saturate the detector. The observations were split in 4 observing blocks, 2 in each filter. The B-band observations were obtained in Dark (fractional lunar illumination, FLI , 30%) time and the I-band in grey time (FLI $\sim$40%). The observations were performed in service mode, under clear conditions and with a seeing from 1 to 1.5 arcsec. The observations used the Standard Resolution (SR) collimator and the red sensitive (MIT) CCD with 2x2 binning offering a pixel scale of 0.25"/pix.
The images were processed using standard techniques, using the ESO/MIDAS software, and using standard calibration data obtained on the same night(s). We first created a master bias which was subtracted from all science and other calibration frames. We then created a normalised flatfield using the twilight calibration frames. A cosmic ray filtering (FILTER/COSMIC) was applied to the science data which were then flat-fielded, corrected for airmass, CCD gain and normalised to 1 s. Residual image artefacts (bad columns and charge bleeding trails) were flagged. Residual sky emission was removed from each frame by fitting a first order polynomial surface (using the routine FIT/FLAT). Finally, all images for each filter were registered and combined. We note that the centre of each knot in Haro 11, as well as some relatively bright neighbouring stars, are saturated in the FORS images. As we here focus on the extended low surface brightness emission around Haro 11, this is not a major concern. For photometric zero-point we used the values for the observing nights from the FORS photometric trending analysis, available on the following ESO webpage: https://eso.org/observing/dfo/quality/FORS2/reports/HEALTH/trend_report_ZEROPOINTS_BH_HC.html
We also verified the calibration for an unsaturated star next to Haro 11 that is also detected in the HST/ACS/WFC/F435W image, available on MAST.